Laser Spectroscopy of the A^{3}Π_{1u} ←X^{1}Σ_{g}^{+} System of I_{2} in the 11650–12300 cm^{−1}region

\underline{TOKIO YUKIYA}, NOBUO NISHIMIYA AND MASAO SUZUKI, Department of Electronic Engineering, Tokyo Institute of Polytechnics, Iiyama 1583, Atsugi City, 243-02 Kanagawa, Japan.

The Doppler limited electronic absorption spectrum in the A^{3}Π_{1u}–X^{1}Σ_{g}^{+} system was measured in the 11650– 12300 cm^{−1}region using a Ti:sapphire ring laser (Coherent 899-21) pumped by an argon ion laser (Coherent Innova 90). The Q-branch lines of J = 10 to 100 belonging to the v′ ← v′′ = (29 ∼ 31)′ ← 0′′, (21 ∼ 30)′ ← 1′′ and (22 ∼ 31)′ ← 2′′ progressions were assigned. The hyperfine splittings of these lines could not be resolved. The wavelength measurements were done using a wavelength-meter (Anritu MF9630A) with an accuracy of ±0.5 ppm. Using a least squares fitting procedure, the spectroscopic constants of ν_{Q}^{0}, B′_{vf} − B′′_{v}, D′_{vf} − D′′_{v} and H′_{v} − H′′_{v} were determined. The standard deviations were smaller than 0.003 cm^{−1}. By using the Dunham’s coefficents of the X-state reported by Tromp and Le Roy (1 ), the spectroscopic constants of T′_{v}, B′_{vf}, D′_{vf} and H′_{vf} suitable for v′ = 21′ ∼ 31′ of the A-state were determined by a global least squares fits. The rotational constant of B′_{vf} at the v′ = 21 agreed with that reported by Ishikawa et al. (2 ) within the standard error, while the discrepancy of 0.1 cm^{−1}was found at the value of T′_{v}.

(1 ) J. W. Tromp and R. J. Le Roy, J. Mol. Spectrosc. 109, 352–367(1985). (2 ) T. Ishikawa, H. Takekawa and K. Obi, J. Mol. Spectrosc. 159, 443–457(1993).